Seven times the sum of twice a number and three is equal to six times the difference of the same number doubled and four. What is the number?
Answer by stanbon(75887)

Seven times the sum of twice a number and three is equal to six times the difference of the same number doubled and four. What is the number?
-------
7(2x+3) = 6(2x-4)
-------
14x + 21 = 12x - 24
-------
2x = -45
-------
x = -22.5
----------
Cheers,
Stan H.
----------